Calculate the area of a rectangle with a length of 13.0 cm and a width of 91.0 mm.
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

13*9.1=118.3 cu cm

Answer by Alan3354(69443)

13 cm = 130 mm
Area = 130*91 = 11830 sq mm